# thanks: Equal contributionthanks: Equal contribution

### v1

@@ III RESULTS
Next, inspired by theoretical models used for modeling RT motion  [47, 48, 49, 50, 51, 52], we quantify this motion in terms of tumbling frequency, denoted by $\lambda$. Visual analysis of trajectories shows that a tumbling event happens when at least one robot undergoes a complete rotation (see SM movie 6). This is always accompanied by $\beta(t)$ peaking at $180^{\circ}$ (Figs. 2(c) and 2(d) for experiment and simulation respectively). Using this criterion, we calculate $\lambda$ as the mean time-frequency of occurrences of such tumble peaks. We find that for $\alpha=90^{\circ}$, $\lambda$ shows a systematic variation as a function of $\delta$ and $D_{r}$, as shown in Fig. 3(a) and 3(b) for experiment and simulation, respectively. Our results show that both $\delta$ and $D_{r}$ promote tumbling in the system along the expected lines as they introduce stochasticity in the run state and increased torque, respectively (Eq. (3b)). The system also shows interesting phase behaviour when we vary the parameter $\alpha$. Experiments and simulation show that the run state becomes considerably stable for $\alpha>90^{\circ}$. This is indicated by regions surrounded by dashed lines corresponding to vanishing $\lambda$ in Fig. 3(c) and (d) for experiment and simulation, respectively. When we plot $\lambda$ as a function of $D_{r}$ for various values of $\alpha$, we find a well-defined critical $D_{r}$ beyond which $\lambda$ increases abruptly indicated by black arrowheads in Figs. 3(e) and 3(f). On the contrary, for $\alpha\leq 90^{\circ}$, $\lambda$ shows a gradual increase with $D_{r}$. We further observe that the mean run speed, $V_{\text{run}}$, calculated by taking the average of $V$ while running, remains independent of $\alpha$ for $\alpha\leq 90^{\circ}$ and decreases with $\alpha$ as $\alpha$ exceeds $90^{\circ}$ (Fig. 3(g) and (h) for experiment and theory respectively). In summary, our system provides various experimental parameters that can be adjusted to modify the tumbling frequency, emphasizing the importance of the physical coupling of active units in the propulsion mechanism.
Figure 4: Understanding RT motion using flow-diagrams of angular coordinates $(\theta_{+},\theta_{-})$: (a) A schematic diagram clarifying the angular coordinates $(\theta_{+},\theta_{-})$. Here, $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$. (b) Flow diagram of $(\theta_{+},\theta_{-})$ for $\alpha=120^{\circ}$ evaluated from the deterministic part $\mathcal{T}_{\pm}$ of Eq.  (8). Blue circles around stable points $S_{1}$ and $S_{2}$ are regions where flow lines converge, representing the run states of the system. Red circles around unstable points $U_{1}$ and $U_{2}$ are regions of diverging flow lines. Experimental trajectories for $D_{r}=0$ are illustrated as black dashed lines (see SM movie 7). Schematic configurations corresponding to $S_{1}$ and $S_{2}$ are shown at the bottom. $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are misaligned by an angle of $2\alpha-\pi=60^{\circ}$ in both cases. (c) Flow diagram for $\alpha=90^{\circ}$. Blue dotted lines at $\theta_{-}=n\pi$ are semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$, where $n$ denotes positive integers. See SM movie 8 for the experimental trajectories of $D_{r}=0$ in black dashed lines. (d) Flow profile for $\alpha=60^{\circ}$. The system exhibits four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, and $U_{4}$). Solid blue lines at $\theta_{-}=n\pi$ represent configurations that are stable in $\theta_{-}$ and neutral in $\theta_{+}$. The black dashed experimental trajectories for the $D_{r}=0$ case closely follow the stable lines (see SM movie 9). The system features infinitely many stable run configurations, corresponding to $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$, a few of which are shown at the bottom. The color bars in all these flow diagrams correspond to the magnitude of the vector $(\mathcal{T}_{+},\mathcal{T}_{-})$ in rad s-1. (e), (f) & (g) show $\theta_{-}$ vs $t$ obtained from the numerical solution of Eq. (8) for $\alpha=120^{\circ}$, 90 and 60, and $D_{r}=0.06$ rad2 s-1. For $\alpha=120^{\circ}$, the plateaus at $\theta_{-}=n\pi-\pi/6$ (indicated by dotted gray lines) correspond to run states near the stable states $S_{1}$ and $S_{2}$. For $\alpha=90^{\circ}$ and $\alpha=60^{\circ}$, the plateaus occur at $\theta_{-}=n\pi$, corresponding to infinitely many run configurations with $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$. Switches between these states occur through configurations where $\mathbf{n_{1}}\nparallel\mathbf{n_{2}}$.
@@
Next, we develop a theoretical framework to explain the observed RT dynamics in our system and to provide a rationale for the $\alpha-D_{R}$ phase diagram. As observed in experiments and numerical simulations, the dynamics of the system at any given time are entirely governed by its internal configuration, which is represented by variables $\theta_{i}$ ($i$ = 1 and 2 for two robots) (Fig. 4(a)). Therefore, we define another set of generalized angular coordinates: $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$ (see Fig. 4(a)). Note that $\beta=|\theta_{\text{p}}|$, where $\theta_{\text{p}}$ is the principal value of the angle $2\theta_{-}$ in the range $[-\pi,\pi]$ and $\theta_{+}$ indicates the average orientation of the robots with respect to the rod. Fig. 4(a) clearly shows that motility causes increasing and decreasing trends in $\theta_{1}$ and $\theta_{2}$, respectively. Therefore, by definition, $\theta_{-}$ also increases with time over large time scales. Moreover, for integers $n$, when $\theta_{-}=n\pi$, the coupled robots move with the maximum possible speed of $v_{a}$. Similarly, when $\theta_{-}=(2n+1)\pi/2$, the system shows tumbling behavior. To elaborate further, we derive the equations of motion for $\theta_{\pm}$ from Eqs. (3b) and (7), using the relation provided in Eq. (5), which are given as follows (also see SM section IIC):
$\dfrac{d\theta_{\pm}}{dt}=\mathcal{T}_{\pm}+\sqrt{2D_{\pm}}\eta_{\pm}(t)\$ (8)
where
$\mathcal{T}_{+}$ $=$ $2v_{a}\sin\theta_{-}\left(\dfrac{1}{\ell}\sin\theta_{+}-\cos\theta_{+}\mathcal{G}\mathcal{H}\right)$ (9)
$\mathcal{T}_{-}$ $=$ $-\dfrac{2\delta v_{a}}{\Gamma_{\tau}}\mathcal{G}\cos^{2}\theta_{+}\cos(\theta_{-}+\alpha)\sin\theta_{-}$ (10)
@@
and $D_{\pm}$ and $\mathcal{H}$ are the functions $\theta_{+}$ , $\theta_{-}$, and $\alpha$ (see SM section IIC). The strengths of the noise terms $D_{\pm}$ in Eq. (8) are of the order of $D_{r}$.
@@
Let us now analyze the effect of the deterministic components $\mathcal{T}_{\pm}$ on the dynamics of $\theta_{\pm}$. In Figs. 4(b)-4(d), we display the flow profiles of the angular coordinates $(\theta_{+},\theta_{-})$ in the absence of noise. The color map represents the magnitude $\mathcal{T}$ in rad s-1 of the vector with two components $(\mathcal{T}_{+},\mathcal{T}_{-})$. Note that the system is invariant under the transformations $\theta_{1}\to\theta_{1}\pm 2\pi$ and $\theta_{2}\to\theta_{2}\pm 2\pi$. Consequently, it remains invariant under $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\pm\pi)$ and $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\mp\pi)$. This implies that the first and third quadrants, as well as the second and fourth, represent the same set of systems. This symmetry is evident in the flow profiles as well.
@@
We begin with the case where $\alpha>90^{\circ}$, with a typical flow profile shown for $\alpha=120^{\circ}$ (Fig. 4(b)). Here, we find two stable and two unstable fixed points denoted by $S_{1}$, $S_{2}$, and $U_{1}$, $U_{2}$ respectively. We also observe an unstable fixed line along $\theta_{-}=n\pi$ which is less relevant to the results presented. Locations of stable points in the first and second quadrants are $(\theta_{+},\theta_{-})$ = $(\pi,3\pi/2-\alpha)$ and $(0,3\pi/2-\alpha)$ for $S_{1}$ and $S_{2}$ respectively, with system configurations shown below Fig. 4(b). Clearly, both configurations lead to a run state. Interestingly, in this case, the stable run states do not correspond to aligned robots (as was the case for $\alpha=90^{\circ}$), but rather to anti-aligned vectors $\mathbf{s}_{i}$. Here, the system favors these stable points for low noise values, resulting in a pronounced run regime at low $D_{r}$ values, as shown in Figs. 3(e) and 3(f). To test this, we conducted experiments for the $D_{r}=0$ case using random initial conditions in $(\theta_{+},\theta_{-})$ and observed that the system follows the predicted flow lines (black dashed lines in Fig. 4(b)) before eventually settling into one of the stable regions, indicated by the dashed circular regions (see SM movie 7). Moreover, as evident from configurations $S_{1}$&\&& $S_{2}$, the system’s run speed at the stable points is given by $v_{a}\sin\alpha$, which decreases with $\alpha$ for $\alpha>90^{\circ}$. This explains the decreasing trend in $V_{\text{run}}$ shown in Figs. 3(g) and 3(h).
An alternate way to visualize RT dynamics is through the time dependence of the parameter $\theta_{-}$. In Fig. 4(e), which corresponds to $D_{r}$ = 0.06 rad2 s-1 in simulation, we find that $\theta_{-}$ always plateaus around $\pi/2-\alpha+n\pi$, which equals $\theta_{-}=n\pi-\pi/6$ for $\alpha=120^{\circ}$ shown by gray dotted lines. However, due to rotational noise, the system occasionally escapes its stable points, causing it to jump between stable states by a value of $\pi$ leading to a tumbling event. As $D_{r}$ increases, the system is more likely to escape the stable points sooner, leading to an increased tumbling rate, as observed in both simulations and experiments.
@@
For $\alpha=90^{\circ}$, no stable fixed points are observed. Instead, there exists a fixed line at $\theta_{-}=n\pi$, which is semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$ (blue dashed line in Fig. 4(c)). Therefore, as the system reaches this line, it runs for a while, before sliding towards increasing $\theta_{-}$ due to noise. Experiments performed at $D_{r}=0$, in this case, show traced paths that agree well with theoretical values (see SM movie 8). Also, as shown in Fig. 4(f), the system still tends to stay closer to $\theta_{-}=n\pi$ (dotted gray lines) due to the neutral nature, although only for a short period, describing the origin of RT dynamics on our system for $\alpha=90^{\circ}$ shown in Fig. 1.
Finally, for $\alpha<90^{\circ}$, and using $\alpha=60^{\circ}$ as a typical example, we find four unstable points and a fixed line at $\theta_{-}=n\pi$ as shown in Fig. 4(d). This fixed line corresponds to infinitely possible configurations where robots are aligned (i.e., $\mathbf{n}_{i}$ are parallel to each other with three typical ones schematically shown below Fig. 4(d)). Consequently, the system favors configurations along this line leading to run states with $V_{\text{run}}\approx v_{a}$ (as observed experimentally in Fig. 3(g)) that become progressively noisier with increasing $D_{r}$. Again, gray dashed line indicates the experimental trajectory for $D_{r}=0$ case (SM movie 9). Furthermore, in Fig. 4(g), we present $\theta_{-}$ as a function of time, which exhibits relatively noisier plateaus corresponding to run states with occasional escapes, equivalent to the behavior observed for $\alpha>90^{\circ}$. A key distinction, however, is that while the system for $\alpha>90^{\circ}$ is limited to two stable run states, an infinite number of run configurations are possible when $\alpha<90^{\circ}$.
Therefore, our theoretical model not only explains the emergence of complex RT motion in experiments but also elucidates how this complex behavior systematically depends on experimental parameters $\alpha$ and $D_{r}$.
IV Conclusion and Discussion
@@
Inspired by microorganism motility, we present a robotic model system that replicates run-and-tumble-like (RT-like) motion. Our minimal experimental setup consists of two self-propelled robots connected by a rigid rod, each exhibiting an overdamped active Brownian (AB) motion despite its mesoscopic, centimeter-scale size. The rigid rod, attached to a pivot point on the robot’s body, rotates freely and facilitates rotational motion. By varying the pivot placement relative to the robot’s polarity axis, we observe a range of rich, tunable RT-like dynamics. We quantify the RT-like motion using the tumbling frequency and identify two key tuning parameters. The first, $\delta$, is the distance of the pivot from the polarity axis, which promotes tumbling frequency by amplifying torque. The second, $\alpha$, is the angle between the pivot-to-center line and the polarity axis, which also adjusts the tumbling frequency based on rotational noise programmed inside each robot. The tunability of RT motion with these parameters highlights the critical role of inter-flagella coupling through the organism’s body in achieving synchronization cycles. We also developed a theoretical model that captures the essential features of our system and reproduces the experimental results, elucidating the intricate dependence of RT dynamics on $\delta$ and $\alpha$.
A distinctive feature of our synthetic system is its ability to mimic overdamped dynamics, a hallmark of microorganisms operating at extremely low Reynolds numbers. Our robots, driven by motorized wheels through mechanical gears and responding to microcontroller signals, undergo rolling-without-slipping motion, thus faithfully emulating overdamped active Brownian motion. We also demonstrate that reducing wheel traction on the substrate by applying oil alters the tumbling event, causing it to lose its characteristic sharp and sudden turns. Moreover, simulation results by numerically solving the theoretical model, which also assumes overdamped rotational and translational dynamics, align well with experimental observations. These results suggest that similar design principles may underlie motility in microscopic organisms.
While our theoretical model qualitatively matches experimental results, it underestimates the tumbling frequency. This might be because of multiple intricate factors in our experimental system. For example, our theoretical model features propulsion force acting along an axis passing through the centre of the robot. Clearly, it ignores the fact that robots are driven through their wheels, and thus, propulsion forces act along two axes passing parallel to them. Additionally, weight distribution and the elevated placement of pivot points introduce additional torques and frictional differences between the wheels, factors omitted in our minimal model. Nevertheless, our theoretical framework uses the general AB model without such complications and, therefore, should capture the essential physics of overdamped dynamics in both living and artificial systems at microscopic scales. Our results thus provide a fundamental understanding of RT-like motion with broad implications for studying motility in active and living systems operating in low Reynolds number regimes.
ACKNOWLEDGEMENTS
@@
NK acknowledges financial support from DST-SERB for CRG grant number CRG/2020/002925 and IITB for the seed grant. HS acknowledges SERB for the SRG (grant no. SRG/2022/000061-G). SP thanks CSIR India for the research fellowship (File no: 09/087(1040)/2020-EMR-I).
Author contribution statement
NK and HS designed the research. SP performed the experiments and analyzed data. UP and HS developed the theoretical model and ran simulations. All authors contributed equally in writing the paper.

### v2

@@ Abstract
Next, inspired by theoretical models used for modeling RT motion  [44, 45, 46, 47, 48, 49], we quantify this motion in terms of tumbling frequency, denoted by $\lambda$ which equals $<\tau_{\text{run}}^{-1}>$. The averaging is performed over all run events. We find that for $\alpha=90^{\circ}$, $\lambda$ shows a systematic variation as a function of $\delta$ and $D_{r}$, as shown in Fig. 3(a) and 3(b) for experiment and simulation, respectively. Our results show that both $D_{r}$ and $\delta$ promote tumbling in the system along the expected lines as they introduce stochasticity in the run state and increased torque, respectively [see Eq. (14)b of SM]. The system also shows an interesting phase behaviour when we vary the parameter $\alpha$. Experiments and simulation show that the run becomes considerably stable for $\alpha>90^{\circ}$. This is indicated by regions surrounded by dashed lines corresponding to vanishing $\lambda$ in Figs. 3(c) and (d) for experiment and simulation, respectively. When we plot $\lambda$ as a function of $D_{r}$ for various values of $\alpha$, we find a well-defined critical $D_{r}$ beyond which $\lambda$ increases abruptly indicated by black arrowheads in Figs. 3(e) and 3(f). On the contrary, for $\alpha\leq 90^{\circ}$, $\lambda$ shows a gradual increase with $D_{r}$. We further observe that the mean run speed, $V_{\text{run}}$, calculated by taking the average of $V$ while running, remains independent of $\alpha$ for $\alpha\leq 90^{\circ}$ and decreases with $\alpha$ as $\alpha$ exceeds $90^{\circ}$ [insets of Fig. 3(e) and (f)].
@@
Now, we perform a theoretical analysis to explain the observed RT dynamics in our system and to provide a rationale for the $\alpha-D_{r}$ phase diagram. As observed in experiments and numerical simulations, the dynamics of the system at any given time are entirely governed by its internal configuration, which is represented by variables $\theta_{i}$ ($i$ = 1 and 2 for two robots) [Fig. 4(a)]. Therefore, we define another set of generalized angular coordinates: $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$ (see Fig. 4(a)). Note that $\beta$ equals the principal value of the angle $2\theta_{-}$ in the range $[-\pi,\pi]$ and $\theta_{+}$ indicates the average orientation of the robots with respect to the rod. Fig. 4(a) clearly shows that motility causes increasing and decreasing trends in $\theta_{1}$ and $\theta_{2}$, respectively. Therefore, by definition, $\theta_{-}$ also increases with time over large time scales. Moreover, for integers $n$, when $\theta_{-}=n\pi$, the robots are parallel to each other and thus move with the maximum possible speed of $v_{a}$. Similarly, when $\theta_{-}=(2n+1)\pi/2$, the system shows tumbling behavior. To elaborate further, we derive the equations of motion for $\theta_{\pm}$ from Eqs. (30) and (32) of SM, which are given as follows (also see SM section IIC):
$\dfrac{d\theta_{\pm}}{dt}=\mathcal{T}_{\pm}+\sqrt{2D_{\pm}}\eta_{\pm}(t),\$ (1)
@@
where $\mathcal{T}_{+}=(2v_{a}/\ell)\sin\theta_{-}\left(\sin\theta_{+}-\ell\cos\theta_{+}\mathcal{G}\mathcal{H}\right)$, $\mathcal{T}_{-}=-(2\delta v_{a}/\Gamma_{\tau})\mathcal{G}\cos^{2}\theta_{+}\cos(\theta_{-}+\alpha)\sin\theta_{-}$, and $D_{\pm}$, $\mathcal{G}$ and $\mathcal{H}$ are the functions $\theta_{+}$ , $\theta_{-}$, and $\alpha$ (see SM section IIC). The strengths of the noise terms $D_{\pm}$ in Eq. (1) are of the order of $D_{r}$. $\eta_{\pm}(t)$ are the delta-correlated noise with zero mean and variance one.
Figure 4: (a) A schematic diagram clarifying the angular coordinates $(\theta_{+},\theta_{-})$. Here, $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$. (b) Flow diagram of $(\theta_{+},\theta_{-})$ for $\alpha=120^{\circ}$ evaluated from the deterministic part $\mathcal{T}_{\pm}$ of Eq.  (1). $S_{1}$\&& $S_{2}$ and $U_{1}$\&& $U_{2}$ encircled by dotted lines represent stable and untable points respectively. Black dashed lines represent experimental trajectories for $D_{r}=0$. Schematic configurations of $S_{1}$ and $S_{2}$ are shown at the bottom. $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are misaligned by an angle of $2\alpha-\pi=60^{\circ}$ in both cases. (c) Flow diagram for $\alpha=90^{\circ}$. Blue dotted lines at $\theta_{-}=n\pi$ are semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$, where $n$ denotes positive integers. Black dashed lines are the experimental trajectories of $D_{r}=0$. (d) Flow profile for $\alpha=60^{\circ}$. The system exhibits four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, and $U_{4}$). Solid blue lines at $\theta_{-}=n\pi$ represent configurations that are stable in $\theta_{-}$ and neutral in $\theta_{+}$. The black dashed experimental trajectories for the $D_{r}=0$ case. Three among infinitely many run configurations corresponding to $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ are shown at the bottom. The color bars in (b), (c) &\&& (d) correspond to the magnitude of the vector $(\mathcal{T}_{+},\mathcal{T}_{-})$ in rad s-1.
@@
Let us now analyze the effect of the deterministic components, $\mathcal{T}_{\pm}$, on the dynamics of $\theta_{\pm}$. In Figs. 4(b)-4(d), we display the flow profiles of the angular coordinates $(\theta_{+},\theta_{-})$ in the absence of noise. The color map represents the magnitude $\mathcal{T}$ in rad s-1 of the vector with two components $(\mathcal{T}_{+},\mathcal{T}_{-})$. Note that the system is invariant under the transformations $\theta_{1}\to\theta_{1}\pm 2\pi$ and $\theta_{2}\to\theta_{2}\pm 2\pi$. Consequently, it remains invariant under $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\pm\pi)$ and $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\mp\pi)$. This implies that the first and third quadrants, as well as the second and fourth, represent the same set of systems. This symmetry is evident in the flow profiles as well.
@@
We begin with the case where $\alpha>90^{\circ}$, with a typical flow profile shown for $\alpha=120^{\circ}$ in Fig. 4(b). Here, we find two stable ($S_{1}$, $S_{2}$) and two unstable ($U_{1}$, $U_{2}$) fixed points. We also observe an unstable fixed line along $\theta_{-}=n\pi$ which is less relevant to the results presented. Locations of stable points in the first and second quadrants are $(\theta_{+},\theta_{-})$ = $(\pi,3\pi/2-\alpha)$ and $(0,3\pi/2-\alpha)$ for $S_{1}$ and $S_{2}$ respectively, with system configurations shown below Fig. 4(b). Both configurations correspond to run states. Therefore, the system favors these stable points for low noise values, resulting in a pronounced run regime at low $D_{r}$ values, explaining the experimental results presented in Figs. 3(c) and (e). Interestingly, both $S_{1}$\&& $S_{2}$ correspond to non-aligned robots ($\mathbf{n_{1}}\nparallel\mathbf{n_{2}}$). To test this, we conduct experiments for the $D_{r}=0$ case using random initial conditions in $(\theta_{+},\theta_{-})$ and find that the system follows the predicted flow lines (black dashed lines in Fig. 4(b)) before eventually settling into one of the stable regions, indicated by the dashed circular regions (see SM movie 6). Moreover, as evident from configurations $S_{1}$&\&& $S_{2}$, the system’s run speed at the stable points is given by $v_{a}\sin\alpha$, which decreases with $\alpha$ for $\alpha>90^{\circ}$. This explains the decreasing trend in $V_{\text{run}}$ shown in insets of Figs. 3(e) and 3(f). We also observe that at low but finite $D_{r}$ values, the system runs corresponding to $\theta_{-}$ = $\pi/2-\alpha+n\pi$, which equals $\theta_{-}=n\pi-\pi/6$ for $\alpha=120^{\circ}$ while occasionally transitioning between $S_{1}$ and $S_{2}$ leading to tumbling events [See SM Section IID].
@@
For $\alpha=90^{\circ}$, no stable fixed points are observed. Instead, there exists a fixed line [blue dashed line in Fig. 4(c)] at $\theta_{-}=n\pi$, which is semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$ Therefore, as the system reaches this line, it runs for a while before sliding towards increasing $\theta_{-}$ due to noise. Experiments performed at $D_{r}=0$, in this case, show traced paths that agree well with theoretical values (see SM movie 7). Finally, for $\alpha<90^{\circ}$, and using $\alpha=60^{\circ}$ as a typical example, we find four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, $U_{4}$) and a fixed line at $\theta_{-}=n\pi$ as shown in Fig. 4(d). This fixed line is neutral in $\theta_{+}$ while stable in $\theta_{-}$ corresponding to infinitely possible configurations with $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ with schematics of three typical configurations shown below the Fig. 4(d). Consequently, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ as observed experimentally in inset of Fig. 3(e). Again, the grey dashed line indicates the experimental trajectory for $D_{r}=0$ (SM movie 8).
@@
To conclude, inspired by microorganism motility, we present a robotic model system that replicates run-and-tumble-like (RT-like) motion. Our robots, driven by motorized wheels through mechanical gears and responding to microcontroller signals, undergo rolling-without-slipping motion. As a result, they faithfully emulate overdamped active motion, a hallmark of microorganisms operating at extremely low Reynolds numbers. We program both robots to exhibit an overdamped active Brownian (AB) motion. The rigid rod, attached to a pivot point on the robot’s body, rotates freely and facilitates rotational motion. By varying the pivot placement relative to the robot’s polarity axis, we observe a range of rich, tunable RT-like dynamics. We quantify the RT-like motion using the tumbling frequency and identify two key tuning parameters. The first, $\delta$, is the distance of the pivot from the robot’s centre, which promotes tumbling frequency by amplifying torque. The second, $\alpha$, is the angle between the pivot-to-center line and the polarity axis, which also adjusts the tumbling frequency based on rotational noise programmed inside each robot. The tunability of RT motion with these parameters might have implications in the critical role played by inter-flagella connection through the organism’s body in achieving synchronization cycles [23, 25]. We also developed a theoretical model that captures the essential features of our system and reproduces the experimental results, elucidating the intricate dependence of RT dynamics on $\delta$ and $\alpha$.
@@
Acknowledgements - NK acknowledges financial support from DST-SERB for CRG grant number CRG/2020/002925 and IITB for the seed grant. HS acknowledges SERB for the SRG (grant no. SRG/2022/000061-G). SP thanks CSIR India for the research fellowship (File no: 09/087(1040)/2020-EMR-I).